Given a normal distribution with a standard deviation of 14, what is the mean if 40% of the values are below 60?


1
Expert's answer
2021-01-19T18:17:32-0500

We have that

"\sigma = 14"

"P(X<60)=0.4"

"P(X<60)=P(Z<\frac{60 - \mu}{\sigma})=0.4"

Z-value of –0.25 corresponds to 40% of the area under the curve.

Then

"\frac{60-\mu}{\sigma}=-0.25 \implies \mu=60+0.25\cdot\sigma=60+0.25\cdot14=63.5"


Answer: the mean is 63.5